# A Gift from the Integration of Discriminative and Diffusion-based Generative Learning: Boundary Refinement Remote Sensing Semantic Segmentation

### v1

@@ III-A Latent Diffusion Model for Generative Semantic Segmentation
During inference, random noise $z_{T}$ is denoised through the denoising model $\epsilon_{\theta}$ via a Markov chain under the guidance of condition $c$, specifically $z_{T}\to z_{T-\Delta}\to\cdots\to z_{0}$, which can be expressed as:
@@
However, latent diffusion models generally assume that sample data is continuous, making them inconvenient for processing discrete label maps. We aim to map discrete label maps to a continuous space while ensuring that the encoded label features can be mapped back to category probability space. Therefore, we additionally learn an embedding network consisting of a shallow encoder $E_{e}$ and a shallow decoder $D_{e}$. $E_{e}$ maps discrete labels to a continuous space, while $D_{e}$ maps encoded label features back to category probability space. The optimization of the embedding network is independent of the diffusion model and does not require image data. Specifically, $E_{e}$ contains an embedding layer that maps discrete labels to a 3-channel continuous space, applies a sigmoid function for normalization, and then scales the encoded labels to the range $[-1,1]$. $D_{e}$ contains two convolutional layers that first map the 3-channel encoded labels to 128 channels, and then to a $K$-channel category probability space, where $K$ is the number of categories. To improve the embedding network’s robustness during training, we add Gaussian noise (variance = 0.25) to $E_{e}$’s output before feeding it to $D_{e}$. We optimize the embedding network by minimizing the cross-entropy loss between the category probabilities produced by $D_{e}$ and the discrete labels.
@@ III-B Rethinking Diffusion Model from a Frequency Domain Perspective for Semantic Segmentation
In the frequency domain, regions composed of similar neighboring pixels in an image typically manifest as low-frequency components, while edge regions with significant differences between adjacent pixels appear as high-frequency components. Similarly, when segmentation maps are transformed into the frequency domain via Fourier transform, the semantic content generally corresponds to low-frequency components, while boundaries between different semantic regions correspond to high-frequency components. Research has shown that diffusion generative models enhance the learning of high-frequency components in image generation tasks [19], [54]. Therefore, we explore whether diffusion models can enhance the refinement of object boundaries in semantic segmentation tasks. In this section, we first qualitatively evaluate the accuracy of diffusion-based generative semantic segmentation in inferring both high-frequency and low-frequency components. Subsequently, we theoretically analyze how the diffusion denoising process effectively enhances the model’s ability to learn high-frequency components from input data.
Figure 3: Analysis of the effectiveness of diffusion-based generative learning for boundary segmentation. (a) Qualitative assessment of inference accuracy on high-frequency and low-frequency components. (b) Theoretical analysis of how the diffusion denoising process enhances high-frequency component learning.
@@ III-B2 Why Diffusion Denoising Process Effectively Enhances High-Frequency Component Learning
If we view the diffusion network as a linear filter, the optimal filter derived from the frequency response at each timestep is commonly known as the Wiener filter. Therefore, some studies have used Wiener filters to analyze the denoising process of diffusion models [54, 55]. In this subsection, we further explain the intrinsic reason why diffusion models effectively enhance high-frequency component learning based on the denoising process of the Wiener filtering method. Assuming $z_{0}$ is a wide-sense stationary signal, and $\varepsilon$ is white noise with variance $\sigma^{2}=1$. For $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\varepsilon$, given a linear denoising Wiener filter $h_{t}$, the objective function of the diffusion model at time $t$ can be redefined as:
@@ III-C1 Coarse Discriminative Semantic Segmentation as Conditional Guidance for Generative Learning
These coarse segmentation results serve as conditional guidance for generative learning, providing large-scale semantic context information to the generative semantic segmentation training process. In diffusion model training, the sample is defined as $z_{0}=E_{S}(y)$, and the conditional information as $c_{I}=E_{I}(x)$ and $c_{r}=E_{S}(y_{r})$. The functions $E_{S}(\cdot)$ and $E_{I}(\cdot)$ denote encoding functions for segmentation maps and remote sensing images, respectively, aimed at transforming input data into low-dimensional latent encoding spaces to reduce computational complexity. The diffusion process transforms sample $z_{0}$ into a noisy sample $z_{t}$. Under the joint conditional guidance of $c_{I}$ and $c_{r}$, we train a Unet model $\epsilon_{\theta}$ to predict the noise added to the noisy sample $z_{t}$.
@@ III-C3 Regularization Strategy through Representation Alignment for Accelerated Training
Diffusion models have garnered significant attention within deep generative modeling due to their robust generation capabilities and strong performance. However, the high computational cost during both training and inference stages presents a substantial obstacle for their practical application in downstream tasks. In the computer vision domain, existing works optimize and accelerate the training process through various strategies, such as leveraging visual foundation models for representation alignment [20],[58] or enhancing semantic associations via masked modeling [59],[60]. Inspired by the REPA method [20], we utilize DINOv2 [61] as the foundational visual model to facilitate joint representation alignment with the diffusion backbone, thereby accelerating training. The detailed procedure is described in the following two steps:
@@
(b) Diffusion Representation Alignment: At each timestep of the diffusion process, the final-layer hidden state from the Unet encoder is denoted by $h_{t}=f_{\theta}(x_{t})$. This hidden features is transformed via a learnable MLP module $h_{\phi}$ to align with $y_{c}$.
@@ III-C4 Training & Inference
        label_latents = vae_encode(embedded_labels)
        rough_latents = vae_encode(embedded_rough)
@@
        preds_noise, latents_repr = model(noisy_latents,
@@
        $\mathcal{L}_{\text{total}}$ = $\mathcal{L}_{\text{MSE}}$(pred_noise, noise) + $\lambda$cdot $\mathcal{L}_{\text{REPA}}$
@@
Label encoding: Diffusion models typically assume continuous input data, conflicting with discrete labels in semantic segmentation. Therefore, we design and pretrain a shallow label encoder $E_{e}$ that maps discrete labels into a continuous differentiable latent space. Specifically, we construct a learnable embedding table for each class, transforming label map $y\in\{0,1,\dots,K-1\}^{H\times W}$ into a continuous embedding feature $E_{0}\in\mathbb{R}^{C\times H\times W}$. This embedding is then normalized to the interval $[-1,1]$ by:
@@
Here, $\mathcal{L}_{\text{MSE}}$ represents the standard denoising mean squared error loss, while $\mathcal{L}_{\text{REPA}}$ is the representation alignment loss. During the first 500 training steps, we set $\lambda=0.5$ to incorporate semantic alignment; subsequently, $\lambda$ is set to 0 to prevent continuous alignment signals from interfering with segmentation learning.
@@
            latents = $\sqrt{1-\bar{\alpha}_{t^{\prime}}}*\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t^{\prime}}}*\hat{\epsilon}$
@@ IV-A Datasets
TABLE I: Remote sensing semantic segmentation data set used in this paper
@@
The CHN6-CUG road dataset aims to address the representation limitations of existing public datasets in complex road scenes. The dataset was collected from 6 typical urban areas and annotated by image interpretation experts to cover 12 road subclasses including railways, highways, urban roads, etc. It contains 4,511 images of 1024×1024 pixels at 0.5-meter resolution. In our experiments, all road subclasses were merged into one class, with non-road areas serving as the background class. The dataset was divided into training, validation, and test sets in an 8:1:1 ratio.
@@
The WHU Building change detection dataset is constructed from aerial images with 0.075-meter resolution from two periods: 2012 (post-disaster) and 2016 (post-reconstruction). Geometric registration at 1.6-pixel accuracy was achieved using 30 ground control points. The original data covers 20.5 square kilometers, with annotated buildings increasing from 12,796 to 16,077, capturing the complete evolution of building clusters after the disaster. In our experiments, we cropped the original images to 512×512 pixels and used an 8:1:1 ratio for training, validation, and test sets.
The Potsdam and Vaihingen datasets originate from the ISPRS Test Project on Urban Classification, 3D Building Reconstruction, and Semantic Labeling. Both datasets provide pixel-wise annotations for impervious surfaces, buildings, low vegetation, trees, cars, and background. The Potsdam dataset contains 38 aerial images with a ground sample distance (GSD) of 5 cm. Each image includes a true orthophoto (TOP) and a corresponding normalized digital surface model (nDSM). The orthophotos are available in three formats: infrared-red-green (IRRG), red-green-blue (RGB), and RGB-infrared composite. Each at 6000 × 6000 resolution. The Vaihingen dataset contains 33 aerial images with a GSD of 9 cm, also including TOP and nDSM data. The TOP images in this dataset are provided in a near-infrared-red-green (NIR-RG) format with a resolution of 2494 × 2064 resolution. All images are cropped to 512 × 512 patches, yielding 3456 training and 2016 test images for Potsdam, and 562 training and 180 test images for Vaihingen.
@@ IV-B1 Implementation Details
To assess IDGBR’s generalization in improving boundary quality of coarse segmentation maps generated by different base models, we conducted further experiments across multiple base segmentation models spanning four architectural paradigms: (1) CNN architectures: DeepLabv3+ [68] and LSKNet [29]; (2) Transformer-based architecture: SegFormer [69]; (3) Self-supervised learning: Using frozen DINOv2 [61] pretrained weights as visual encoder with Mask2Former’s fine-tunable segmentation head; (4) Generative model architecture: DDP [52] with end-to-end predictions. Coarse labels from the first three paradigms served as conditional inputs for IDGBR boundary optimization, while DDP predictions served only as a baseline for comparing segmentation results.
For the base discriminative training, all base discriminative segmentation models used default hyperparameters and were implemented using open-source libraries. We employed standard data augmentation strategies: (1) random resizing with aspect ratio preserved between 0.5 and 2.0 (output size 512×512); (2) random cropping at size 512 ensuring no class exceeds 75% coverage; (3) horizontal flipping with 50% probability; (4) random adjustments to brightness, contrast, and saturation. All models were trained with a batch size of 8 for 20k iterations.
During the training of IDGBR, we employ a Unet network initialized with the pretrained weights of Stable Diffusion v1.5 as the backbone for the denoising process. Both images and labels are separately encoded using VAE encoders, and the resulting representations are concatenated to serve as conditional inputs. To avoid biases arising from model configuration, the training and inference procedures of IDGBR are kept consistent across different tasks and backbone models; for different numbers of classes, only the output projection dimension is modified during inference. Specifically, IDGBR model is trained for 80,000 iterations with a batch size of 4 and a fixed learning rate of 1e-5. Inference is performed over 25 steps, with a classifier-free guidance weight of 3. To accelerate model convergence, a representation alignment mechanism is introduced during early training. Concretely, feature maps from the final layer of the Unet downsampling stage are first extracted and projected to semantic dimensions using a three-layer linear projection module, while a single linear expansion module handles spatial dimensional matching. The transformed features are then aligned with the semantic representations extracted by DINOv2, using cosine similarity as the alignment objective. This mechanism is only activated during the first 200 training steps to guide the model toward more efficient semantic learning.
@@ IV-C Evaluation on Binary Segmentation Tasks
In binary Segmentation, we construct a validation framework using three representative binary remote sensing tasks: road segmentation, building extraction, and cropland delineation. These three tasks correspond to testing ground object features with linear extension, geometric regularity, and texture complexity, respectively, effectively evaluate IDGBR’s boundary reconstruction capability and target perception capabilities for objects with different morphological characteristics across various discriminative baseline models.
@@ IV-C1 Evaluation on Road Segmentation
The complex topological structures and multi-scale characteristics of roads in the CHN6-CUG dataset provide an effective evaluation benchmark for evaluating the model’s capability in boundary refinement and connectivity restoration capabilities for linear features. Comparative results are presented as follows:
TABLE II: Quantitative results (%) of different methods on the CHN6-CUG road dataset.
Method DiscModel Backbone Road Kappa OA WFm (3px B)
IoU F1score
DDP ConvNext-t 64.19 78.19 77.76 97.59 36.52
DeepLabV3+ ResNet-50 62.59 76.99 75.66 97.48 34.44
@@
As shown in Table II, analysis based on the WFm boundary metric demonstrates that optimizing boundaries with the proposed IDGBR framework leads to significant improvements in boundary scores for models such as DeepLabV3+, SegFormer, DINOv2, and LSK, indicating strong cross-model generalization capability. In terms of the IoU metric, all baseline models achieved consistent improvements after boundary optimization, with DINOv2 achieving the most notable gains. In addition, although the DDP model, as a diffusion-based benchmark, achieves overall accuracy comparable to discriminative methods after adaptation to downstream tasks, its WFm remains lower than that of models optimized by IDGBR.
@@ IV-C2 Evaluation on Cultivated Land Segmentation
TABLE III: Quantitative results (%) of different methods on FGFD.
Method DiscModel Backbone Farmland Kappa OA WFm (3px B)
IoU F1score
DDP ConvNext-t 87.44 93.30 83.86 92.16 42.40
DeepLabV3+ ResNet-50 86.15 92.56 82.56 91.47 39.66
@@
As shown in Table III, IDGBR achieved an average WFm improvement exceeding 9.08%, significantly improving the usability of extracted field parcels. The IDGBR-refined segmentation results of DINOv2 showed the greatest enhancement, with WFm rising 13.13% to 49.78%. For other models, IDGBR boosted WFm by 9.69% on DeepLabV3+, 7.94% on SegFormer, and 5.56% on LSK, further demonstrating its effectiveness.

### v2

@@ III-A Latent Diffusion Model for Generative Semantic Segmentation
During inference, random noise $z_{T}$ is denoised through the denoising model $\epsilon_{\theta}$ via a Markov chain under the guidance of condition $c$, specifically $z_{T}\to z_{T-1}\to\cdots\to z_{0}$, which can be expressed as:
@@
However, latent diffusion models generally assume that sample data is continuous, making them unsuitable for processing discrete label maps. To address this, we introduce an auxiliary embedding network consisting of a shallow encoder $E_{e}$ and a decoder $D_{e}$. Specifically, $E_{e}$ projects discrete labels into a continuous embedding space to facilitate diffusion modeling, while $D_{e}$ maps the encoded features back to the category probability space. The optimization of the embedding network is independent of the diffusion model and does not require image data.
@@ III-B Rethinking Diffusion Model from a Frequency Domain Perspective for Semantic Segmentation
In the frequency domain, regions composed of similar neighboring pixels in an image typically manifest as low-frequency components, while edge regions with significant differences between adjacent pixels appear as high-frequency components. Similarly, when segmentation maps are transformed into the frequency domain via Fourier transform, the semantic content generally corresponds to low-frequency components, while boundaries between different semantic regions correspond to high-frequency components. Research has shown that diffusion generative models enhance the learning of high-frequency components in image generation tasks [18, 49]. Therefore, we explore whether diffusion models can enhance the refinement of object boundaries in semantic segmentation tasks. In this section, we first qualitatively evaluate the accuracy of diffusion-based generative semantic segmentation in inferring both high-frequency and low-frequency components. Subsequently, we theoretically analyze how the diffusion denoising process effectively enhances the model’s ability to learn high-frequency components from input data.
Figure 3: Analysis of the effectiveness of diffusion-based generative learning for boundary segmentation. (a) Qualitative assessment of inference accuracy on high-frequency and low-frequency components. (b) Theoretical analysis of how the diffusion denoising process enhances high-frequency component learning. The horizontal axis represents relative spatial frequency ($f$), where smaller $|f|$ values correspond to low-frequency structures and larger $|f|$ values represent high-frequency details; the vertical axis is the normalized frequency response. The curves illustrate the dynamic evolution of the retention of different frequency components during the reverse denoising process.
@@ III-B2 Why Diffusion Denoising Process Effectively Enhances High-Frequency Component Learning
If we view the diffusion network as a linear filter, the optimal filter derived from the frequency response at each timestep is commonly known as the Wiener filter. Therefore, some studies have used Wiener filters to analyze the denoising process of diffusion models [49, 50]. In this subsection, we further explain the intrinsic reason why diffusion models effectively enhance high-frequency component learning based on the denoising process of the Wiener filtering method. Let $z_{0}$ denote a wide-sense stationary signal and $\varepsilon$ represent white noise with variance $\sigma^{2}=1$. For $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\varepsilon$, given a linear denoising Wiener filter $h_{t}$, the objective function of the diffusion model at time $t$ can be redefined as:
@@ III-C1 Coarse Discriminative Semantic Segmentation as Conditional Guidance for Generative Learning
These coarse segmentation results serve as conditional guidance for generative learning, providing large-scale semantic context information to the generative semantic segmentation training process. In diffusion model training, the sample is defined as $z_{0}=E_{S}(y)$, and the conditional information as $c_{I}=E_{I}(x)$ and $c_{r}=E_{S}(y_{r})$. The functions $E_{S}(\cdot)$ and $E_{I}(\cdot)$ denote encoding functions for segmentation maps and remote sensing images, respectively, aimed at transforming input data into low-dimensional latent encoding spaces to reduce computational complexity. The diffusion process transforms sample $z_{0}$ into a noisy sample $z_{t}$. Under the joint conditional guidance of $c_{I}$ and $c_{r}$, we train a Unet model $\epsilon_{\theta}$ to predict the noise contained in the noisy sample $z_{t}$.
@@ III-C3 Regularization Strategy through Representation Alignment for Accelerated Training
Although diffusion models excel at learning high-frequency patterns, their substantial training costs limit practical application. Inspired by REPA [19], we propose a regularization strategy based on representation alignment during early training stages. For semantic segmentation, this serves two key purposes: (1) it stabilizes early training by providing a robust semantic prior that guides optimization, leading to faster convergence; and (2) it enforces feature consistency across varying appearances and contexts, encouraging the model to reason based on semantic structure rather than superficial local textures. To this end, we select DINOv2 [53] as the semantic anchor, leveraging its strong general semantic representation—learned via self-distillation on large-scale unlabeled data—which focuses on scene geometry and object structure rather than low-level textures, thereby providing unbiased semantic guidance and avoiding task-specific bias. The implementation details are as follows:
@@
(b) Diffusion Representation Alignment: At each timestep of the diffusion process, the final-layer hidden state from the Unet encoder is denoted by $h_{t}=f_{\theta}(x_{t})$. This hidden feature is transformed via a learnable MLP module $h_{\phi}$ to align with $y_{c}$.
@@ III-C4 Training & Inference
     label_latents = vae_encode(labels_embedded)
     rough_latents = vae_encode(rough_embedded)
@@
[…]
@@
     $\mathcal{L}_{\text{REPA}}$ = $\mathbb{E}\!\bigl[\text{cosine\_similarity}(\text{latents\_repr},\,\text{imgs\_repr})\bigr]$
@@
Label encoding: Diffusion models typically assume continuous input data, conflicting with discrete labels in semantic segmentation. Therefore, we design and pretrain a shallow label encoder $E_{e}$ that maps discrete labels into a continuous differentiable latent space. Specifically, we construct a learnable embedding table for each class, transforming label map $y\in\{0,1,\dots,K-1\}^{H\times W}$ into a continuous embedding feature $E_{0}\in\mathbb{R}^{C\times H\times W}$. This embedding is then normalized to the interval [1,1][-1,1] by:
@@
Here, $\mathcal{L}_{\text{MSE}}$ represents the standard denoising mean squared error loss, while $\mathcal{L}_{\text{REPA}}$ is the representation alignment loss. During the first 200 training steps, we set $\lambda=0.5$ to incorporate semantic alignment; subsequently, $\lambda$ is set to 0 to prevent continuous alignment signals from interfering with segmentation learning.
@@
      latents = $\sqrt{\bar{\alpha}_{t^{\prime}}}*\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t^{\prime}}}*\hat{\epsilon}$
@@ IV-A Datasets
TABLE I: Remote sensing semantic segmentation datasets used in this paper
@@
The CHN6-CUG road dataset aims to address the representation limitations of existing public datasets in complex road scenes. The dataset was collected from 6 typical urban areas and annotated by image interpretation experts to cover 12 road subclasses including railways, highways, urban roads, etc. It contains 4,511 images of 512 × 512 pixels at 1-meter resolution. In our experiments, all road subclasses were merged into one class, with non-road areas serving as the background class. The dataset was divided into training and test sets in a 4:1 ratio.
@@
The WHU Building change detection dataset is constructed from aerial images downsampled to 0.3-meter resolution from two periods: 2012 (post-disaster) and 2016 (post-reconstruction). Geometric registration at 1.6-pixel accuracy was achieved using 30 ground control points. The original data covers 20.5 square kilometers, with annotated buildings increasing from 12,796 to 16,077, capturing the complete evolution of building clusters after the disaster. In our experiments, we cropped the images to 512 × 512 pixels, utilizing the post-disaster data for training and testing (7:3 ratio) and the post-reconstruction data exclusively for testing.
The Potsdam and Vaihingen datasets originate from the ISPRS Test Project on Urban Classification, 3D Building Reconstruction, and Semantic Labeling. Both datasets provide pixel-wise annotations for impervious surfaces, buildings, low vegetation, trees, cars, and background. The Potsdam dataset contains 38 aerial images with a ground sample distance (GSD) of 5 cm. Each image includes a true orthophoto (TOP) and a corresponding normalized digital surface model (nDSM). The orthophotos are available in three formats: infrared-red-green (IRRG), red-green-blue (RGB), and RGB-infrared composite, each at 6000 × 6000 resolution. The Vaihingen dataset contains 33 aerial images with a GSD of 9 cm, also including TOP and nDSM data. The TOP images in this dataset are provided in a near-infrared-red-green (NIR-RG) format with a resolution of 2494 × 2064. All images are cropped to 512 × 512 patches, yielding 3456 training and 2016 test images for Potsdam, and 562 training and 180 test images for Vaihingen.
@@ IV-B1 Implementation Details
To provide a comprehensive evaluation of the generalization capability of IDGBR, we selected representative general segmentation methods covering three mainstream paradigms. By utilizing these models to generate the diverse coarse semantic priors required for boundary optimization, we verify the framework’s robustness across varying initial inputs: (1) Convolutional Neural Network methods: DeepLabV3+ [60] and the large-kernel CNN LSKNet [25] optimized for remote sensing tasks; (2) Transformer methods: SegFormer [61]; (3) Self-supervised foundation models: using frozen DINOv2 [53] weights as the visual encoder.
For the baseline discriminative training, all models were implemented using open-source libraries with default hyperparameters. We employed standard data augmentation strategies: (1) random resizing with aspect ratio preserved between 0.5 and 2.0 (output size 512 × 512); (2) random cropping at size 512 ensuring no class exceeds 75% coverage; (3) horizontal flipping with 50% probability; (4) random adjustments to brightness, contrast, and saturation. All models were trained with a batch size of 8 for 20k iterations.
During the training of IDGBR, we employ a Unet network initialized with the pretrained weights of Stable Diffusion v1.5 as the backbone for the denoising process. Both images and labels are separately encoded using VAE encoders, and the resulting representations are concatenated to serve as conditional inputs. To avoid biases arising from model configuration, the training and inference procedures of IDGBR are kept consistent across different tasks and backbone models; for different numbers of classes, only the output projection dimension is modified during inference. Specifically, IDGBR model is trained for 80,000 iterations with a batch size of 4 and a fixed learning rate of 1e-5. Inference is performed over 25 steps, with a classifier-free guidance weight of 3. To accelerate model convergence, a representation alignment mechanism is introduced during early training. Concretely, we extract feature maps from the final layer of the Unet downsampling stage. These features are projected to semantic dimensions using a three-layer linear projection module, while a single linear expansion module handles spatial dimensional matching. The transformed features are then aligned with the semantic representations extracted by DINOv2, using cosine similarity as the alignment objective. This mechanism is only activated during the first 200 training steps to guide the model toward more efficient semantic learning.
@@ IV-C Evaluation on Binary Segmentation Tasks
In binary segmentation, we construct a validation framework using three representative binary remote sensing tasks: road segmentation, building extraction, and cropland delineation. These three tasks correspond to testing ground object features with linear extension, geometric regularity, and texture complexity, respectively, and effectively evaluate IDGBR’s boundary reconstruction capability and target perception capabilities for objects with different morphological characteristics across various discriminative baseline models.
@@ IV-C1 Evaluation on Road Segmentation
The complex topological structures and multi-scale characteristics of roads in the CHN6-CUG dataset serve as a rigorous benchmark for assessing the boundary refinement and connectivity restoration for linear features. Comparative results are presented as follows:
TABLE II: Quantitative comparison (%) of our proposed method against various general segmentation methods on the CHN6-CUG road dataset.
Method DiscModel Backbone Road Acc WFm (3px B)
IoU F1-score
DeepLabV3+ ResNet-50 62.59 76.99 97.48 34.44
@@
As shown in Table II, analysis based on the WFm boundary metric demonstrates that optimizing boundaries with the proposed IDGBR framework leads to significant improvements in boundary scores for models such as DeepLabV3+, SegFormer, DINOv2, and LSK, indicating strong cross-model generalization capability. In terms of the IoU metric, all baseline models achieved consistent improvements after boundary optimization, with DINOv2 achieving the most notable gains.
@@ IV-C2 Evaluation on Cultivated Land Segmentation
TABLE III: Quantitative comparison (%) of our proposed method against various general segmentation methods on the FGFD.
Method DiscModel Backbone Farmland Acc WFm (3px B)
IoU F1-score
DeepLabV3+ ResNet-50 86.15 92.56 91.47 39.66
@@
As presented in Table III, IDGBR significantly enhances the usability of extracted field parcels, notably achieving a maximum WFm improvement of 13.13% with the DINOv2 baseline (rising from 33.60% to 46.73%). For other architectures, IDGBR consistently boosted performance, increasing WFm by 9.69% on DeepLabV3+, 7.94% on SegFormer, and 5.56% on LSK, further demonstrating its robust effectiveness.